can the product rule be used to verify the chain rule? support your answer with example.


1
Expert's answer
2022-01-20T12:32:56-0500

"\text{Yes, for a single variable}\\
\text{Let x, y be two given functions}\\ \text{Consider} \, f=(x+y)^2. \\
\text{By the chain rule, we have}\\
f'=2(x + y)(x' + y') = 2(xx'+ xy'+ yx'+ yy')\\
\text{Now, expand f to obtain} \, f=x^2+2xy + y^2\\
\text{Thus, we have}\\
f' =2xx'+2(xy)' +2yy' = 2(xx'+(xy)'+yy'))\\
\text{It follows immediately that}\\
(xy)'=xy' + yx' \\
\text{Thus, we see that equality holds}"